Generating the index and chage log for the Babel system

Johannes Braams

1994/06/05

1 Description

This file contains the source for the style files for the MakeIndex program. These are needed to produce the index for the complete source listing and to produce the listing of the changes.

2 The index and the list of changes

Most of the entries are the same as in the files gind.ist and gglo.ist from the doc package.

1 ⟨∗idx | glo⟩

 2 actual '='
 3 quote '!'
 4 level '>'

5 ⟨/idx | glo⟩

Because the babel system consists of so many files the default codeline numbering scheme of the doc package has been adapted. The line numbers consist of two parts seperated with a dot. This has to made known to the MakeIndex program when it produces the index.

 6 ⟨+idx⟩page_compositor "."

Any further entries are again copied from gind.ist and gglo.ist.

7 ⟨∗idx⟩

 8 preamble
 9 "\n \\begin{theindex} \n \\makeatletter\\scan@allowedfalse\n"
10 postamble
11 "\n\n \\end{theindex}\n"

12 ⟨/idx⟩ 13 ⟨∗glo⟩

14 preamble
15 "\n \\begin{theglossary} \n
16  \\makeatletter\\scan@allowedfalse\n"
17 postamble
18 "\n\n \\end{theglossary}\n"
19 keyword "\\glossaryentry"

20 ⟨/glo⟩

21 ⟨+idx | glo⟩item_x1   "\\efill \n \\subitem "

22 ⟨+idx⟩item_x2   "\\efill \n \\subsubitem "
23 ⟨+glo⟩item_x2   "\\ "

24 ⟨∗idx | glo⟩

25 delim_0   "\\pfill "
26 delim_1   "\\pfill "
27 delim_2   "\\pfill "
28 % The next lines will produce some warnings when
29 % running Makeindex as they try to cover two different
30 % versions of the program:
31 lethead_prefix   "{\\bfseries\\hfil "
32 lethead_suffix   "\\hfil}\\nopagebreak\n"

33 ⟨/idx | glo⟩

34 ⟨+idx⟩lethead_flag       1
35 ⟨+glo⟩lethead_flag       0

36 ⟨∗idx | glo⟩

37 heading_prefix   "{\\bfseries\\hfil "
38 heading_suffix   "\\hfil}\\nopagebreak\n"

39 ⟨/idx | glo⟩

40 ⟨+idx⟩headings_flag       1
41 ⟨+glo⟩headings_flag       0